OpenStudy (anonymous):

which of the numbers below could be terms in the sequence an=4n-7 A)29 B)21 C)20 D)28

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{a_n=4n-7\\~\\ \color{red}{option~~A.)}29=4n-7\\~\\ 4n=36\\~\\ n=9\\~\\ \text{as this is integer it can be a possible term}\\~\\ \color{red}{option~~B.)}21=4n-7\\~\\ 4n=28\\~\\ n=7\\~\\ \text{as this is integer it can be a possible term}\\~\\ \color{red}{option~~C.)}20=4n-7\\~\\ 4n=27\\~\\ n=\dfrac{27}{4}\\~\\ \text{as this is not integer it cant be a possible term}\\~\\ \color{red}{option~~D.)}28=4n-7\\~\\ 4n=35\\~\\ n=\dfrac{35}{4}\\~\\ \text{as this is not integer it cant be a possible term}\\~\\}\end{align}\)

OpenStudy (anonymous):

so A & B is the answer

OpenStudy (mathmath333):

\(\Huge\tt \bf \begin{align} \color{black}{\checkmark}\end{align}\)

OpenStudy (anonymous):

thank you!